OpenStudy (anonymous):

P=80-Q/2 turn the equation around so that Q is a function of P (that is, add q/2 to both sides and subtract p from both, then solve for q). I got Q=160-2P

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

I want the answer like this P= Q

OpenStudy (mertsj):

You have it. Q= 160-2P

OpenStudy (anonymous):

but I want the answer P=

OpenStudy (anonymous):

2P=160-Q?

OpenStudy (anonymous):

I want to get rid of 2

OpenStudy (mertsj):

That's the way the equation was given: P=80-Q/2

OpenStudy (anonymous):

2P=160-Q. Would it be work?

OpenStudy (mertsj):

Are you talking about finding the inverse: switching P and Q and then solving for P again?

OpenStudy (anonymous):

No Can I switch 2p and Q?

OpenStudy (mertsj):

No. You can have the given equation in a variety of forms. It is given where P is a function of Q. If you rearrange it so that it says Q= then Q is a function of P

OpenStudy (anonymous):

I See. thank you!

OpenStudy (mertsj):

yw